The domain of r(t)=⟨sin(t),cos(t),ln(t+1)⟩ is
Select one:
a. R

b. (−1,∞)

c. (−∞,−1]


d. (−π,2π)


e. (π,2π)

f. (−∞,−2]
1
Expert's answer
2020-05-18T20:06:17-0400

"\sin t : t\in(-\infty, +\infty),\\
\cos t : t\in(-\infty, +\infty),\\
\ln(t+1): t+1>0 \to t>-1\to t\in(-1,+\infty)" .

The correct answer is b.